OpenStudy (anonymous):

i thought i was doing this one right but the answer i got wasn't one of my choices. so im asking for help. Solve: 3^(x)=(1)/(81)

OpenStudy (saifoo.khan):

\[3^x = \frac{1}{81} \] \[x = \log_3 81^{-1}\] \[x = \log_3 (3^{4})^{-1}\] \[x = \log_3 3^{-4}\] \[x = -4 \log_3 3\]Can you do from here?

hero (hero):

Or you can do \(3^x = 3^{-4}\) and conclude that x = -4

hero (hero):

@saifoo.khan

OpenStudy (saifoo.khan):

Oh right. that's a more shortcut.

OpenStudy (anonymous):

thanks both of you :)